### v1

# Thermodynamic and quantum fluctuations of horizon area

Abstract
The event horizon is a source of irreversibility, analogous to statistical irreversibility. This is why for systems with an event horizon there is no difference between quantum and thermal fluctuations. Quantum processes of quantum tunneling determine the thermodynamics of these systems, their temperatures, entropies and fluctuations. We considered three examples of entropy variance that support this point of view: (i) the variance of the area of the black hole horizon, obtained by consideration of quantum fluctuations; (ii) the variance of the entropy of the Hubble volume in the de Sitter state, obtained by consideration of thermal fluctuations; and (iii) the variance of entropy in integers in the Planckon model, determined by the Poisson distribution.
pacs:
I Introduction
We discuss the quantum and thermal fluctuations of the entropy related to the event horizons. In Section II the quantum fluctuations of the black hole area are considered in terms of the pair of the canonically conjugate gravitational variables. In Section III the thermal fluctuations of the entropy of the Hubble volume in the de Sitter state are considered. The variance of the thermal entropy exactly corresponds to the quantum fluctuations of the area of the horizon. This demonstrates the connection between quantum and thermodynamic fluctuations in the systems with event horizon. In Section IV the thermodynamic and quantum fluctuations of black hole are discussed in the toy model, which is suggested by the non-extensive Tsallis-Cirto statistics. It is ensemble of the correlated pairs of Planckons, where each Planckon is the object with reduced Planck mass and zero entropy.
II Quantum fluctuations of black hole entropy
Here we consider quantum fluctuations of the area of the black hole horizon using the gravitational canonically conjugate variables. In the case of the black hole, these are the horizon area $A$ and the gravitational coupling $K=1/(4G)$. These canonically conjugate variables were used in particular for the calculations of the transition rate from the black hole to the white hole with the same mass.Volovik2022 This rate is determined by the integral $\int_{C}A(K)dK$ over the tunneling trajectory $C$.
For the canonically conjugate variables one has the standard uncertainty relation ($\hbar=c=1$):
$\Delta A\,\Delta K\geq\frac{1}{2}\,$, (1)
The horizon area is $A=\pi M^{2}/K^{2}$, and thus for fixed mass $M$ of the black hole, the variation of the gravitational coupling leads to the following variation the horizon area:
$\frac{\Delta A}{A}=2\frac{\Delta K}{K}\,.$ (2)
Inserting variation $\Delta K=(K/2A)\Delta A$ to Eq.(1) one obtains the area fluctuations:
$\left<(\Delta A)^{2}\right>\,\geq\frac{1}{K}\left<A\right>=4G\left<A\right$>\,. (3)
The coefficient $4$ in the area variance in Eq.(3) differs from the corresponding coefficients in Refs. Ciambelli2025 ; ParikhPereira2024 . However, the same equation (3) can be obtained from the thermodynamic fluctuations of the horizon entropy in the de Sitter (dS) spacetime, see next Section.
III Thermodynamic fluctuations of de Sitter horizon entropy
As distinct from the black hole the dS spacetime is homogeneous and isotropic, and thus can be described by the local thermodynamics.Volovik2025b This thermodynamics has the local entropy density:
$s=3KH=3\pi KT\,.$ (4)
Here $H$ is the Hubble parameter and $T=H/\pi$ is the local temperature of the de Sitter state. This local temperature is twice the Gibbons-Hawking temperature $T_{\rm GH}=H/2\pi$ associated with the cosmological horizon. Nevertheless, the total entropy $S=sV_{H}$ of the Hubble volume $V_{H}=(4\pi/3)/H^{3}$ coincides with the Gibbons-Hawking entropy of cosmological horizon, $sV_{H}=A/4G$. This demonstrates, that the local entropy obeys the holographic bulk-horizon correspondence.Hooft1993 ; Susskind1995 This holographic correspondence connects the extensive entropy of the homogeneous Universe with the non-extensive entropy of the cosmological horizon.
The factor 2 has a natural explanation from the point of view of two observers. The observer measuring the local temperature, such as the activation temperature of the ionization of atoms in de Sitter environment, has all the necessary information. On the other hand, the observer measuring the temperature of the Hawking radiation from the cosmological horizon has no information about the simultaneous creation of the Hawking partner beyond the horizon. This is the reason why this observer underestimates the temperature by a factor of 2. These arguments do not apply to a black hole, where the creation of a Hawking partner is represented by a back reaction.Wilczek2000
Since the de Sitter state is homogeneous and isotropic, we can apply the standard equations for the thermodynamic fluctuations (see the book Landau_Lifshitz by Landau and Lifshitz). Then using the linear dependence of local entropy on temperature and thus the linear dependence of the heat capacity, one obtains the fluctuations of the total entropy in the Hubble volume $V_{H}$:
$\left<(\Delta S)^{2}\right>=\left<S\right$>\,. (5)
Since $S=KA=A/4G$, this gives again the Eq.(3). Note that Eq.(5) does not contain the gravitational coupling $K$. This is in agreement with the observation by Jacobson Jacobson1994 that both $K$ and $S$ are renormalized by the very same quantum fluctuations.
The exact relation between the quantum and thermodynamic fluctuations is typical for the systems with horizons. This can be seen also from the processes of splitting of the black hole into the smaller parts. On one hand, such process is quantum, being determined by the macroscopic quantum tunneling, and on the other hand it is determined by the decrease of entropy after splitting,Wilczek2000 ; Volovik2022 which is manifestation of the thermal fluctuations.
[…]
where $\left<{\cal H}\right>=A/4G$, see e.g. Ref. VerlindeZurek2020 and the references in Ref. Ciambelli2025 .
IV Black hole fluctuations via ensemble of Planck scale ”atoms”
Another interesting property of the quantum-thermodynamic fluctuations of the black hole is provided by the toy model of the black hole entropy.Volovik2025P In this model, the black hole with mass $M$ is represented by the peculiar ensemble of $N$ Planckons – objects with reduced Planck massMarkov1967 ; Hawking1971 ; Aharonov1987 ; Rovelli2024 ; YenChinOng2025 ; Trivedi2025 :
$M=Nm_{\rm P}\,\,,\,\,m_{\rm P}^{2}=\frac{K}{2\pi}=\frac{1}{8\pi G}\,.$ (7)
The processes of splitting and merging of the black holes demonstrate that the entropy $S_{\rm BH}(N)$ of the black hole is not determined by the individual Planckons, which have zero entropy, $S(N=1)=0$. The entropy of black hole is determined by the correlated (or entangled) pairs of Planckons. Each pair has entropy $S(N=2)=1$, and the entropy of the black hole is determined by the number $\cal N$} of pairs:
$S_{\rm BH}(N)={\cal N}=\begin{pmatrix}N\\ 2\end{pmatrix}=\frac{N!}{2!(N-2)!}=\frac{N(N-1)}{2}\,.$ (8)
In the thermodynamic limit $N\gg 1$, the equation (5) gives for fluctuations of the Planckon number:
$\left<(\Delta N)^{2}\right>=\frac{1}{2}\,.$ (9)
Such unusual variance of the number of Planckons $N$, which is not proportional to $N$, is the consequence of the Tsallis-Cirto non-extensive statisticsTsallisCirto2013 ; Tsallis2020 with $\delta=2$. This statistics describes the non-extensive composition law for entropy in the processes of splitting and merging of black holes: (see Refs.Volovik2025TC ; Manoharan2025 and references therein):
$\sqrt{S(N_{1}+N_{2})}=\sqrt{S(N_{1})}+\sqrt{S(N_{2})}\,$, (10)
This non-extensive composition law is in the origin of the black-hole thermodynamics.
On the other hand, the ensemble of the $\cal N$} pairs of Planckons has conventional composition law $S({\cal N}_{1}+{\cal N}_{2})=S({\cal N}_{1})+S({\cal N}_{2})$ and conventional fluctuations, which are consistent with the Poisson distribution:
$\left<(\Delta{\cal N})^{2}\right>={\cal N}\,.$ (11)
This is the representation of the equations (3), (5) and (6) for quantum and thermodynamic fluctuations in integer numbers. In principle, this may suggest the realization of the Bekenstein idea of quantization of the horizon entropy.Bekenstein1974
V Discussion
The event horizon is the source of irreversibility, which is similar to the statistical irreversibility.Witten2025 That is why for systems with horizon, there is no difference between the quantum and thermal fluctuations. The quantum processes of quantum tunneling determine the thermodynamics of these systems, with their temperatures, entropies and fluctuations. We considered three examples of the entropy variance, which support this view.
(i) The variance of the area of the black hole horizon can be obtained in the domain of quantum fluctuations. Here we used the uncertainty principle, which relates the variances of the canonically conjugate variables: the gravitational coupling $K=1/4G$ and black hole area $A$. This gives Eq.(3) for the variance of the horizon area.
(ii) The variance of the entropy of the cosmological horizon in de Sitter state can be obtained in the domain of classical thermodynamics fluctuations. The entropy of the horizon is equal to the entropy of the Hubble volume, thus representing the holographic bulk-horizon connection. As a result, the variance of the entropy can be obtained from the conventional thermodynamics in bulk, which determines thermal fluctiations of thermodynamic variables, such as entropy. The result obtained in the domain of thermal fluctuations coincides with the result obtained for black hole in the domain of quantum fluctuations in Eq.(3).
(iii) The variance of entropy can be expressed in terms of integer numbers $N$ in the Planckon model, which simulates the Tsallis-Cirto non-extensive ensemble with $\delta=2$. The variance is determined by the Poisson distribution of the correlated pairs of Planck scale objects – Planckons. For large number of Planckons the result agrees with the results obtained in (i) and (ii). The Planckon model simulates the Bekenstein-type quantization of the horizon entropy.